### v2

# Robust nonlinear state-feedback control of second-order systems

Abstract
This note proposes a novel nonlinear state feedback controller for perturbed second-order systems. In analogy to a linear proportional-derivative (PD) feedback control, the proposed nonlinear scheme uses the output state of interest and its time derivative, while achieving robust control in a finite time. The control has only one free design parameter, and the closed-loop system is shown to be uniformly asymptotically stable in the presence of matched disturbances. We derive a strict Lyapunov function for the closed-loop control system with a bounded exogenous perturbation, and use it for both the control parameter tuning and analysis of the finite-time convergence. Apart from the numerical results, a revealing experimental example is also shown in favor of the proposed control and in comparison with PD and sub-optimal nonlinear damping regulators.
I Introduction
Convergence properties of a transient response and the ability to compensate for unknown (external) disturbances are often the trade-offs when designing a feedback control at large. Well known, a PD regulator, which is equivalent to the state-feedback control for second-order systems, is sufficient for the unperturbed case provided that an exponential asymptotic convergence is adequate. When the system behavior is subject also to static errors, an integral control action can be added, thus resulting in a standard PID (or integral-plus-state feedback) scheme in case of the linear feedback regulation [1]. From another side, an integral feedback action increases inherently the system order and can also lead to the wind-up effects and even destabilization [2]. If disturbances are not static and possess certain regularity only (e.g., Lipschitz continuity or/and boundedness), then the robust feedback regulators like, for instance, high-gain (see, e.g., [3] and references therein) or finite-time (see, e.g., [4] and a recent text [5]) controllers can be required. Beyond that, the input-to-state stability (ISS) theory [6] becomes necessary for studying the stability and robustness in presence of external disturbances, once the overall feedback system is nonlinear [7].
Second-order systems capture a relatively large class of the (perturbed) dynamic processes, especially in mechanical, electro-magnetic, and fluid-flow domains. In that way, a perturbed double-integrator represents an often encountered system process. A comparative overview of different feedback control principles applied to a classical double-integrator can be found in [8]. Also the sliding-mode control methods, see [9, 10], become relatively popular for compensating the unknown matched perturbations. Indeed, a feedback control action proportional to the sign of the control error has already been recognized in earlier works as particularly efficient for a time-optimal stabilization [11, 12], and disturbance compensation in perturbed systems [13]. However, a discontinuous control signal, especially a continuously switching one, can be undesirable for several controlled processes, especially with regard to the wear effects, overloading of actuators, noise, and energy consumption. Therefore, a feedback control law that would allow both, a sign driven robustness to the unknown but bounded perturbations and such continuous damping action which would minimize the steady-state oscillations, can be desirable. The idea of an inverse to the output’s norm damping in the second-order systems was originally proposed in the so-called sub-optimal nonlinear damping control [14]. Recently, the sub-optimal nonlinear damping control was augmented by a linear integral term in [15], with additional analysis, so as to cope with the matched constant disturbances. Also the ISS property was established for time-varying bounded perturbations in that augmented control system [15].
The present work builds up on the idea proposed in [14] and uses the sign-feedback principle of sliding mode control theory, while introducing a novel nonlinear state feedback control for the perturbed second-order systems. The unknown time-varying perturbations are assumed to be upper bounded, and the controller has only one free design parameter, similarly to [14]. The introduced nonlinear state feedback controller has the same structure as its linear counterpart (i.e., the PD one) and feeds back the output of interest and its time derivative. Moreover, it offers finite-time convergence despite unknown bounded perturbations and is continuous during convergence to the globally stable equilibrium at the origin. Particularly remarkable in comparison to several robust second-order sliding mode controllers is that the proposed control offers convergence without overshooting the origin (or reference value) that comes in favor of numerous potential applications.
The note is organized as follows. The proposed nonlinear state feedback control is introduced in Section II, also demonstrating qualitatively the phase portrait and shape of the converging state trajectories and control value, and that in comparison with a linear state feedback control (equivalent to PD) and sub-optimal nonlinear damping control [14]. The main part of the control system analysis is delivered in Section III. First, the system is shown to be uniformly asymptotically stable, then the finite-time convergence is formally proved. Section IV reports the experimental evaluation of the proposed control and compares it with an equivalent PD and sub-optimal nonlinear damping controllers. The results reported in this communication are briefly summarized in Section V.
II Novel nonlinear feedback control
In this section, first, we rigorously describe the studied control problem; second, the proposed control algorithm is presented; third, we finish with an exemplary illustration of the behavior for the closed-loop dynamics.
II-A Problem statement
We consider a sufficiently general class of perturbed second-order systems:
$\frac{\mathrm{d}^{2}y(t)}{\mathrm{d}t^{2}}=u(t)+d(t),\quad t\geq 0,$ (1)
with the control signal $u(t)\in\mathbb{R}$. The matched unknown disturbance quantity $d(t)\in\mathbb{R}$ is assumed to be Lebesgue measurable and essentially bounded, where $\|d\|_{\infty}=\text{ess}\sup_{t\geq 0}|d(t)|\leq D$ with a known upper bound $D>0$. Both dynamic system states, $x(t)=\left(x_{1}(t),x_{2}(t)\right)^{\top}\equiv\bigl{(}y(t),\dot{y}(t)\bigr{)}^{\top}\in\mathbb{R}^{2}$ (including output derivative [^1] We will use the doted symbol, i.e., $\dot{a}\equiv\mathrm{d}a/\mathrm{d}t$, to denote the time derivative of a variable $a$, equally as for the first order differential equations.), are available for a feedback design, while the robust finite-time stabilization at $x=0$ is our primary goal.
More precisely, it is required to design a state feedback $u(t)=u\bigl{(}x_{1}(t),x_{2}(t)\bigr{)}$ guaranteeing that there exist $\sigma,T\in\mathcal{K}$ such that $|x(t)|\leq\sigma(|x(0)|)$ for all $t\geq 0$ and $|x(t)|=0$ for all $t\geq T(|x(0)|)$ for any $x(0)\in\mathbb{X}=\{x\in\mathbb{R}^{2}\,:\,x_{1}\neq 0\}\cup\{0\}$ and $\|d\|_{\infty}\leq D$. Thus, we assume that at the initial time instant, the position $x_{1}(0)$ is not zero, and it is required to bring $x(t)$ to zero in a finite time independently in the presence of a properly bounded disturbance $d$.
Note that in case of an additional (inertial) scaling constant in the left hand-side of (1), like in the experimental study of a moving lumped mass shown in Section IV, this should be additionally incorporated as a normalization factor of the control terms, cf. [16].
II-B Control law
The proposed second-order nonlinear state-feedback control
$u(t)=\begin{cases}-|x_{1}(t)|^{-1}\bigl{(}\gamma x_{1}(t)+|x_{2}(t)|x_{2}(t)\bigr{)}&x_{1}(t)\neq 0\\ -\gamma\,\text{sign}\bigl{(}x_{1}(t)\bigr{)}&x_{1}(t)=0\end{cases},$ (2)
where $\gamma>D$ is the single control design parameter, yields for $x(t)\in\mathbb{X}\setminus\{0\}$ the closed-loop system dynamics to
$\frac{\mathrm{d}x_{1}(t)}{\mathrm{d}t}=x_{2}(t),$ (3)
$\frac{\mathrm{d}x_{2}(t)}{\mathrm{d}t}=\qquad\;\;$
$\begin{cases}-|x_{1}(t)|^{-1}\bigl{(}\gamma x_{1}(t)+|x_{2}(t)|x_{2}(t)\bigr{)}+d(t)&x_{1}(t)\neq 0\\ -\gamma|x_{1}(t)|^{-1}x_{1}(t)+d(t)&x_{1}(t)=0\end{cases}.$
As we can conclude, the proposed control is discontinuous at $x_{1}=0$, and in the set $\mathbb{X}$ the irregularity happens at the origin only, which is the equilibrium of the system. To be more precise, there are two discontinuous terms. The first one is the conventional sign function $x_{1}(t)|x_{1}(t)|^{-1}$ widely used in the sliding mode control theory. Hence, for definition of solutions in the face of this discontinuity, the Filippov’s theory can be used with convex embedding [9, 10]. The second discontinuity can come from the fraction $|x_{2}(t)|x_{2}(t)|x_{1}(t)|^{-1}$, which requires a special attention at $x_{1}=0$, so that another analysis approach will be utilized in the following. Below, in order to substantiate existence and uniqueness of solutions for (3), we will demonstrate similarity of the system (3) via a time change to another system with well defined solutions, as we have already done for the sub-optimal nonlinear damping control in [15]. Next, the stability and performance of the closed-loop system can be analyzed. Before doing it, we proceed with some illustrative simulation results showing the intuition behind the proposed design and qualitative properties of the state trajectories of the controlled system.
II-C Exemplary
An illustration of phase-portrait of the closed-loop control system (3) is given in Fig. 1. Here the control gain factor $\gamma=100$ is arbitrary assigned. Due to a symmetric behavior with respect to the origin, the trajectories for different initial conditions are shown in the 1st and 4th quadrants only.
Figure 1: Phase portrait exemplary of the closed-loop control system (3).
The trajectories of $x_{2}(t)$ and the control values $u(t)$ are also qualitatively compared in Fig. 2 between the proposed control (3), sub-optimal nonlinear damping control [14], and critically damped linear state feedback control.
Figure 2: Response of the dynamic state $x_{2}(t)$ in (a) and control signal $u$ in (b) for the proposed control (3), sub-optimal nonlinear damping control [14], and critically damped linear state feedback control.
All three controllers are assigned with the same output feedback gain factor $\gamma=100$, while the response of $x_{2}(t)$ is depicted on the logarithmic scale for the sake of a better comparison, see Fig. 2 (a). The corresponding control values $u(t)$ are shown in Fig. 2 (b). All three feedback controllers disclose the same control value at the beginning, this is due to the equal gaining factor $\gamma$ and the same initial condition $x_{1}(0)=-1$, $x_{2}(0)=0$. Then, one can recognize that the control (3) uses the range of admissible values more efficiently during the convergence phase.
III Analysis
We start this section with proof of existence and uniqueness of solutions in (3). Next, the uniform asymptotic stability at the origin will be substantiated. The rate of convergence is evaluated at the end of the section.
III-A Existence of solutions
In the Appendix, for the system (9) operating in time $\tau\geq 0$ the following properties are demonstrated for any $x(0)\in\mathbb{R}^{2}$ and the inputs with $\|d\|_{\infty}\leq D$:
  1. 1.
    global existence and uniqueness of solutions for all $t\geq 0$,
  2. 2.
    invariance of the sets $\mathcal{X}=\{x\in\mathbb{R}^{2}:x_{1}=0\}$ and $\mathbb{X}=\left(\mathbb{R}^{2}\setminus\mathcal{X}\right)\cup\{0\}$,
  3. 3.
    existence of the equilibrium at the origin, which is uniformly globally asymptotically stable.
To use these facts in the analysis of (3), introduce a new (scaled) time argument for (9):
$t:=\phi(\tau)=\int\limits_{0}^{\tau}\bigl{|}x_{1}(s)\bigr{|}ds,$ (4)
which is well-defined, and $t\in\mathbb{R}_{+}$ for $\tau\in\mathbb{R}_{+}$ (indeed, $t=0$ for $\tau=0$, and $\lim_{\tau\to+\infty}\int\limits_{0}^{\tau}\bigl{|}x_{1}(s)\bigr{|}ds\leq+\infty$). Moreover, ϕ:++:italic-$\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a monotonously growing function for $x(0)\in\mathbb{X}\setminus\{0\}$ since the origin in such a case is reached for $\tau\to+\infty$ only, and it has an inverse $\tau=\phi^{-1}(t)$. However, the domain of values of $t$ may be bounded since $\bigl{|}x_{1}(\tau)\bigr{|}\to 0$ as $\tau\to+\infty$. That implies $\mathrm{d}t=\bigl{|}x_{1}(\tau)\bigr{|}\mathrm{d}\tau$ while transforming (9) to the closed-loop system (3) for $x(0)\in\mathbb{X}\setminus\{0\}$. Indeed, with a slight ambiguity define $x(t)=x(\phi^{-1}(t))=x(\tau)$ and calculate its derivative in the new time $t$ keeping in mind that $d/dt\,\phi^{-1}(t)=\Bigl{|}x_{1}\bigl{(}\phi^{-1}(t)\bigr{)}\Bigr{|}^{-1}$, which gives that $x(t)$ is a solution of (3):
$\frac{dx(t)}{dt}=\Bigl{.}\frac{dx(\tau)}{d\tau}\Bigr{|}_{\tau=\phi^{-1}(t)}\frac{d\phi^{-1}(t)}{dt}.$
Therefore, the systems (9) and (3) have the same solutions in $\mathbb{X}$ (the origin is the equilibrium for both systems), where the only difference is that each solution becomes scaled along the independent time variable based on the norm of $x_{1}$.
Since (9) has a unique solution asymptotically approaching the origin for any initial condition $x(0)\in\mathbb{X}$ and input with $\|d\|_{\infty}\leq D$, the system (3) also admits this solution while demonstrating the same behavior in finite or infinite time. In the former case, since the origin is the equilibrium of (3), the zero solution can be extended for all $t\geq 0$. Therefore, the following result has been proven:
Proposition 1.
The system (3) admits an unique solution for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ well defined with $x(t)\in\mathbb{X}$ for all $t\geq 0$.
In fact, via such a similarity of (9) and (3) through the time scaling (4), a stronger property is substantiated: uniform asymptotic stability of the origin for (3) in $\mathbb{X}$. However, below we would like to apply the Lyapunov function method directly to (3), so as to establish this property with demonstration of the finite-time convergence rate.
III-B Uniform asymptotic stability
We consider first an unperturbed closed-loop dynamics (3), i.e., with $d=0$. For that system, we next suggest the non-strict Lyapunov function candidate
$E(x)=\gamma|x_{1}|+\frac{1}{2}\,x_{2}^{2},$ (5)
which is positive definite for all $x\in\mathbb{R}^{2}$. Evaluating the time derivative of (5) along the solutions of (3) we obtain
$\dot{E}=\gamma\,\mathrm{sign}(x_{1})\dot{x}_{1}+x_{2}\dot{x}_{2},$
and substituting the right hand-side of (3) into $\dot{E}$ results in
$\dot{E}=-|x_{2}|x_{2}^{2}\leq 0,$
which is negative semi-definite only. It is enough to guarantee boundedness of the state trajectories $x(t)$ in the closed-loop system for all $t\geq 0$.
Now, we are in the position to prove the uniform asymptotic stability of the origin for the perturbed system (3) with $d\neq 0$.
Theorem 2.
The origin for the system (3) is uniformly asymptotically stable for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ if
$\gamma>D^{1.5}+D+\frac{1}{2}.$ (6)
Proof:
As in the Appendix for the system (9), consider for (3) a strict Lyapunov function candidate
$V(x)=\frac{1}{2}\,\left(\begin{array}[]{c}z_{1}\\ x_{2}\end{array}\right)^{\top}\,\left(\begin{array}[]{cc}2\,\gamma&\varepsilon\\ \varepsilon&1\end{array}\right)\left(\begin{array}[]{c}z_{1}\\ x_{2}\end{array}\right)$ (7)
with $z_{1}=\sqrt{|x_{1}|}\,\mathrm{sign}(x_{1})$, which is positive definite for $\sqrt{2\gamma}>\varepsilon$. This Lyapunov function is continuously differentiable for $x\in\mathbb{X}\setminus\{0\}$. Taking the time derivative of $V$ on this set one obtains
$\dot{V}$ $=$ $x_{2}\,d-x_{2}^{2}\frac{|x_{2}|}{|x_{1}|}-\varepsilon\bigl{(}\gamma-\mathrm{sign}(x_{1})\,d\bigr{)}\,\sqrt{|x_{1}|}$
$+\,\varepsilon\frac{x_{2}^{2}}{\sqrt{|x_{1}|}}\,\Bigl{(}0.5-\mathrm{sign}(x_{1})\mathrm{sign}(x_{2})\Bigr{)}.$
Using the Young’s inequality [17] which postulates
$ab\leq\frac{1}{p}\,a^{p}+\frac{p-1}{p}\,b^{p/(p-1)}$
for any $a,b\in\mathbb{R}_{+}$ and $p>1$, one can show that
$\dot{V}\leq-\biggl{(}\frac{2}{3}-\varepsilon\biggr{)}\,x_{2}^{2}\frac{|x_{2}|}{|x_{1}|}-\biggl{[}\varepsilon\Bigl{(}\gamma-\frac{1}{2}-D\Bigr{)}-\frac{2}{3}D^{1.5}\biggr{]}\,\sqrt{|x_{1}|}$ (8)
for any $x\in\mathbb{X}$ and $|d|\leq D$. It can be seen, that taking
$\frac{2}{3}\frac{D^{1.5}}{\gamma-\frac{1}{2}-D}<\varepsilon<\frac{2}{3}$
both expressions in the brackets on the right hand-side of $\dot{V}$ are positive under (6). Hence, $\dot{V}<0$ for $V\neq 0$ in $\mathbb{X}$ that completes the proof. ∎
The low bound of the control gain $\gamma$, which is satisfying the parametric condition (6), is shown as a function of the maximal disturbance magnitude, i.e., $D$, in Fig. 3.
Figure 3: Low bound of control gain depending on perturbation bound.
III-C Finite-time convergence
In order to prove the finite-time convergence (see, e.g., [4] for details) of the system (3), which can be supposed from the numerical results given in Fig. 2 (a), we make further use of the strict Lyapunov function $V$ given in (7).
Theorem 3.
The system (3) is finite-time convergent to the origin for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ provided that (6) is satisfied.
Proof:
The estimate (8) implies boundedness of solutions of (3) in $\mathbb{X}$ for the inputs with magnitude smaller than $D$, with their asymptotic convergence to the origin. Let us consider an auxiliary variable $\zeta(t)=x_{2}(t)x_{1}(t)^{-1}$, which is well defined on the solutions of the system (3) (in the set $\mathbb{X}\setminus\{0\}$), and it has the dynamics:
$\dot{\zeta}=\frac{d\,x_{1}-\bigl{(}\gamma|x_{1}|+\text{sign}(x_{1})|x_{2}|x_{2}\bigr{)}-x_{2}^{2}}{x_{1}^{2}}$
$=\frac{d\,\text{sign}(x_{1})-\gamma}{|x_{1}|}-|\zeta|\zeta-\zeta^{2}$
$=\frac{d\,\text{sign}(x_{1})-\gamma}{|x_{1}|}-\begin{cases}2\zeta^{2}&\zeta>0\\ 0&\zeta\leq 0\end{cases}.$
Note that $(d\,\text{sign}(x_{1})-\gamma)|x_{1}|^{-1}<0$ due to (6), which implies that for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ there is $t_{1}\geq 0$ such that $\zeta(t)<0$ for all $t\geq t_{1}$, i.e., the trajectories enter in the second or the fourth quadrants after finite transients, where the variable $\zeta$ stays separated with zero. Next, consider another auxiliary variable $z(t)=x_{2}^{2}(t)|x_{1}(t)|^{-1}$, which admits the dynamics:
$\dot{z}=\frac{|x_{2}|}{|x_{1}|}\,\Bigl{(}2\text{sign}(x_{2})\bigl{(}d-\gamma\,\text{sign}(x_{1})\bigr{)}$
$-\bigl{(}2+\text{sign}(x_{1})\text{sign}(x_{2})\bigr{)}z\Bigr{)}.$
For $t\geq t_{1}$ and taking into account (6), $\text{sign}(x_{2})(d-\gamma\text{sign}(x_{1}))>0$ out the origin (the lines $x_{1}=0$ and $x_{2}=0$ can be reached only there). Hence, for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$, there is $t_{2}\geq 0$ such that $z(t)\geq\theta>0$ for all $t\geq t_{2}$ (while the trajectories do not converge to zero), where $\theta$ depends only on the choice of $\gamma$. So, there are two possibilities, either the trajectories approach the origin in a finite time, or the variable $z(t)$ becomes strictly positive and separated with zero, which returning to the Lyapunov function $V(x)$ from (7) and the estimate (8) implies for $t\geq t_{2}$:
$\dot{V}\leq-\biggl{(}\frac{2}{3}-\varepsilon\biggr{)}\,\theta|x_{2}|-\biggl{[}\varepsilon\Bigl{(}\gamma-\frac{1}{2}-D\Bigr{)}-\frac{2}{3}D^{1.5}\biggr{]}\,\sqrt{|x_{1}|}$
$\leq-\kappa\sqrt{V}$
for some $\kappa>0$. The latter estimate implies a finite-time convergence of $V$ to zero, according to [4]. ∎
III-D Simulations
The controlled output behavior of (3) is exemplary visualized in Fig. 4 (a). Here a periodic disturbance $d(t)=50\sin(20\pi t)$ is used while the control gain is assigned as $\gamma=100$. Both the disturbance signal and the control value are shown opposite to each other in Fig. 4 (b).
Figure 4: Output response of the perturbed control system (3) in (a), and the control signal versus disturbance value in (b).
IV Experimental control results
The proposed control (3) is experimentally evaluated on the electro-magneto-mechanical actuator system with one translational degree of freedom, see Fig. 5. The system dynamics can be reduced to the model (1).
Figure 5: Experimental setup of the actuator system.
Technical details, including the identified system model and most relevant hardware parameters of the setup, can be found in [16]. The sampling rate is $10$ kHz. We note that the overall upper-bounded perturbation $d$ is due to the constant gravity term, but also due to the time-varying friction and cogging force by-effects of the electro-magneto-mechanical actuator.
The step response of the control (3) (denoted further as NLG due to the nonlinear feedback gains) is compared once with a standard proportional-derivative (PD) control and once with the sub-optimal nonlinear damping control [14] (denoted further as SOND).
The benchmarking linear PD control
$u(t)_{\mathrm{PD}}=-\gamma x_{1}-\sigma x_{2},$
is designed first so that the unperturbed closed-loop control system has two critically damped real poles, i.e., with an optimal value of $\sigma$ that corresponds to the unity damping ratio. This control is denoted by $\mathrm{PD}_{1}$. Then, its output feedback gain $\gamma$ is enhanced while allowing also for the conjugate complex pole pair with a corresponding transient overshoot. That PD-control realization is denoted by $\mathrm{PD}_{2}$.
The parameters of all four evaluated controllers are listed in Table I.
TABLE I: Evaluated controllers
param. \  cont. $\mathrm{PD}_{1}$ $\mathrm{PD}_{2}$ $\mathrm{SOND}$ $\mathrm{NLG}$
$x_{1}$-gain $\gamma$ $500$ $750$ $750$ $25$
$x_{2}$-gain $\sigma$ $2$ $4$- --
For a fair comparison, the SOND control is assigned with the same gaining factor $\gamma=750$ as the faster PD-control $\mathrm{PD}_{2}$. The feedback gain of the NLG-control $\gamma=25$ is intentionally set an order of magnitude lower than the feedback gains of the PD and SOND controllers. This is for demonstrating that the NLG-control can behave less aggressive during the transient response and, at the same time, achieve a much higher accuracy at steady-state. All experimentally measured control responses, to the same reference step of $y_{\mathrm{ref}}=8$ mm, are shown opposite each other in Fig. 6.
Figure 6: Measured step response of PD, SOND, and NLG controllers.
From the response of the NLG-control, one can also recognize a nearly uniform behavior of $x_{2}(t)$ during the whole transient phase, cf. Fig. 2 (a). Note that the corresponding inclination of the $x_{1}(t)$ slope depends on the assigned $\gamma$ parameter. It is also worth noting that a residual non-zero, however minor, steady-state error of the NLG-control is due to the sensor’s bias and noise, while the latter is additionally impacting the used $x_{2}(t)$ signal, which is obtained via differentiation and filtering of the measured $x_{1}(t)$. The experimentally obtained signal $x_{2}(t)$ (when using the NLG controller) is also shown in Fig. 7 for the sake of clarity.
Figure 7: The velocity state $x_{2}(t)$ of NLG control experiment.
V Summary
A novel robust nonlinear state-feedback control for second-order systems was introduced. The control law combines the ideas of the state-dependent nonlinear damping introduced in [14] and the sliding mode control theory, which equips the overall feedback law with an output-sign-dependent term able to compensate for bounded perturbations. The resulted control has only one tuning parameter, the overall output feedback gain, whose selection should be based on the upper bound of the perturbation magnitude, and this closed analytic dependence was shown in the paper. We provided a detailed analysis of uniform asymptotic stability property of the control system and finite-time convergence based on the proposed strict Lyapunov function. The control signal appears to be continuous during the transients, apart from the convergence instant, where the state value reaches the origin (similarly to some high order sliding mode control algorithms). Remarkable features of the proposed control are that: (i) it is equivalent (from the structural viewpoint) to a standard linear PD feedback control and, at the same time, (ii) it can guarantee (without additional integral action) for full compensation of the bounded matched disturbances in the second-order systems without overshoot, while (iii) providing a global finite-time convergence. The proposed control was already evaluated experimentally on an electro-magneto-mechanical actuator system with considerable bounded disturbances. Analysis of its robustness abilities against process and measurement noise and development of the methods for its discretization can be considered as directions of future research.